OpenStudy (anonymous):

Given the literal equation xy+x/z=2+z, with y = 4x + 2, what is the value of z if x = 1?

OpenStudy (anonymous):

plug y in: x(4x+2) + x/z = 2+z plug 1 in for x: 4(1)^2 + 2(1) + 1/z = 2+ z 6 + 1/z = 2 + z z - 1/z = 4 multiply both sides by z: z^2 -1 = 4z \[z ^{2} - 4z -1 = 0\] use quadratic equation. \[\frac{ 4\pm \sqrt{(-4)^{2}-4(1)(-1)} }{ 2(1) }\] z = the resulting 2 terms.